OpenStudy (anonymous):

Help please! what is the soultion radical 5x+1 - radical x=5 x = 0 x = 16 and x = 0 x = 16 x = 16 and x = 1

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{5x+1}-\sqrt{x}=5\] \[\large \sqrt{5x+1}=5+\sqrt{x}\] \[\large \left(\sqrt{5x+1}\right)^2=\left(5+\sqrt{x}\right)^2\] \[\large 5x+1=\left(5+\sqrt{x}\right)^2\] \[\large 5x+1=\left(5+\sqrt{x}\right)\left(5+\sqrt{x}\right)\] \[\large 5x+1=5\left(5+\sqrt{x}\right)+\sqrt{x}\left(5+\sqrt{x}\right)\] \[\large 5x+1=25+5\sqrt{x}+\sqrt{x}*5+\sqrt{x}*\sqrt{x}\] \[\large 5x+1=25+5\sqrt{x}+5\sqrt{x}+\left(\sqrt{x}\right)^2\] \[\large 5x+1=25+10\sqrt{x}+x\] \[\large 5x+1-x-25=10\sqrt{x}\] \[\large 4x-24=10\sqrt{x}\] \[\large \left(4x-24\right)^2=\left(10\sqrt{x}\right)^2\] \[\large 16x^2-192x+576=100x\] \[\large 16x^2-192x+576-100x=0\] \[\large 16x^2-292x+576=0\] I'll let you finish. Make sure to check any/all solutions you get.

OpenStudy (jhannybean):

Thanks jim!!

jimthompson5910 (jim_thompson5910):

sure thing